OpenStudy (anonymous):

guys, let's solve :D \[\int\limits_{}(\sec^2xdx)/\sqrt{4-tg^2x}\]

OpenStudy (angela210793):

solve this urself :P:P i'm too busy reading jokes and chatting :P:P do not disturb :P

OpenStudy (anonymous):

:/

OpenStudy (nottim):

Me too...

OpenStudy (anonymous):

You could use trig identities.

OpenStudy (angela210793):

Einstein....wht i wrote earlier means: i don't know how to solve this :'( and i feel so ignorant :(:( sorry

OpenStudy (anonymous):

:( okay :D

OpenStudy (slaaibak):

i would start by setting tg^2 = k so you get sec^x/sqrt(4-kx) Just looks nicer.

OpenStudy (anonymous):

k(x) you mean? it's a function

OpenStudy (anonymous):

Do you remember this identity?:\[\sec^2x=1+\tan^2x\]

OpenStudy (anonymous):

yeah I remember

OpenStudy (anonymous):

1+tan^2(x) / sqrt(4−tg^2(x))

OpenStudy (anonymous):

* dx

OpenStudy (anonymous):

Mmm it look even worst.

OpenStudy (anonymous):

Yeah,, Can't we make it look like integral of tan^m(x)sec^n(x)dx it would be easier

OpenStudy (slaaibak):

Why is tg^2 a function of x?

OpenStudy (anonymous):

Lets try u = tan x, so du = sec^2xdx

OpenStudy (anonymous):

tan(x), tan^2(x) tan is a function so you can't say that tan^2 is a constant like k, like 2, you can't send it out the integral... because it's a fuction, tan(x)

OpenStudy (anonymous):

\[\int\limits\frac{du}{\sqrt{4-u^2} }\]

OpenStudy (anonymous):

yahh, but we should get rid of this squareroot

OpenStudy (slaaibak):

I never said tan(x) is a constant

OpenStudy (anonymous):

but you tried to do t * x

OpenStudy (anonymous):

you took tan as a number, or a variable

OpenStudy (slaaibak):

Wait. I never took tan as a number. are you saying t=tan?

OpenStudy (anonymous):

you wrote kx => k * x ==> tan^2 * x

OpenStudy (slaaibak):

Where does the tan(x) come from?

OpenStudy (anonymous):

tan is a function...

OpenStudy (anonymous):

that's why it has an y value for every other x values. beucase it's y = tan(x)

OpenStudy (slaaibak):

what is g? and what is t?

OpenStudy (turingtest):

I was treating them as constants

OpenStudy (anonymous):

tg= tan

OpenStudy (earthcitizen):

yh, use substitution method

OpenStudy (slaaibak):

... are you kidding me? How should we know tg = tan? I thought they were constants

OpenStudy (turingtest):

what do you mean? so what is tg^2?

OpenStudy (anonymous):

x.O tg= tan, it sometimes used like that

OpenStudy (anonymous):

tg^2(x) == tan^2(x)

OpenStudy (turingtest):

ok then :/ starting over...

OpenStudy (anonymous):

@slaaiba, Its very common to use tan instead of tg

OpenStudy (slaaibak):

woah, I've never heard of that notation in my life :/ guess you learn something new everyday

OpenStudy (slaaibak):

It should be a fairly simple trig substitution then

OpenStudy (anonymous):

Yahh, I generally learn, and as No-data said, it's very common to use tg

OpenStudy (anonymous):

TANgent line.

OpenStudy (anonymous):

right, you just forgot the dx

OpenStudy (slaaibak):

Can this work: u = tan x du = sec^2 x dx so it's integral of du/sqrt(4-u^2)

OpenStudy (turingtest):

\[\int\frac{\sec^2xdx}{\sqrt{4-\tan^2x}}\]

OpenStudy (anonymous):

yahh

OpenStudy (turingtest):

sure does :)

OpenStudy (anonymous):

it's what No-data did slaaibak

OpenStudy (slaaibak):

Then use the arcsine function

OpenStudy (slaaibak):

Oh, I didn't see, mybad!

OpenStudy (anonymous):

No, it's not that easy slaaibak

OpenStudy (turingtest):

trig sub u=2sin(theta), yes?

OpenStudy (slaaibak):

Isn't \[\int\limits {1\over \sqrt{4-u^2}} du = \arcsin({u \over 2}) + C ?\]

OpenStudy (mr.math):

Yes @slaaibak.

OpenStudy (anonymous):

yes arcsin(tgx/2) + C , right answer but how? :D

OpenStudy (slaaibak):

So then arcsin(tanx/2) + C

OpenStudy (mr.math):

The integral \(\int \frac{1}{\sqrt{1-u^2}} du=\sin^{-1}u+c.\)

OpenStudy (anonymous):

@slaaibak Sorry, you were right

OpenStudy (anonymous):

Ohhh, now I got it

OpenStudy (mr.math):

If you don't know that formula, and you want to derive it yourself, you can use a trig substitution \(u=2\sin z\), and it will lead you to the same result.

OpenStudy (mr.math):

as TurningT just said.

OpenStudy (anonymous):

Ok :D

OpenStudy (turingtest):

\[\int\frac{du}{\sqrt{4-u^2}}\]\[u=2\sin\theta\to du=2\cos\theta d\theta\]\[\int\frac{\cos\theta}{\sqrt{1-\sin^2\theta}}=\int\sec\theta d\theta= \ln|\sec \theta+\tan \theta|+C\]then run the whole thing in reverse to get back to x

OpenStudy (turingtest):

or wait, I messed up...

OpenStudy (slaaibak):

should have been 1, not sec(theta) below the line it would have been cos(theta), not cos^2(theta)

OpenStudy (turingtest):

yeah, oops, I forgot to take the square root on the bottom...

OpenStudy (turingtest):

just integral d(theta)

OpenStudy (anonymous):

\[\int\limits d\theta\]

OpenStudy (slaaibak):

yeah

OpenStudy (mr.math):

Yep!:D \[=\int d\theta=..\]

OpenStudy (turingtest):

\[\int\frac{du}{\sqrt{4-u^2}}\]\[u=2\sin\theta\to du=2\cos\theta d\theta\]\[\int\frac{\cos\theta d\theta}{\sqrt{1-\sin^2\theta}}=\int d\theta =\theta+C\]then run the whole thing in reverse to get back to x

OpenStudy (turingtest):

I just like practicing the Latex...

OpenStudy (slaaibak):

That was typed quite fast

OpenStudy (mr.math):

I think he copied pasted somehow :P

OpenStudy (turingtest):

I saved it just in case I messed up, and just had to change parts...

OpenStudy (mr.math):

I knew it! :P

OpenStudy (slaaibak):

haha, well played mrmath

OpenStudy (mr.math):

lol! I should start doing that. Sometimes, I miss up and have to start all over again!

OpenStudy (anonymous):

Yeah, before my pc formatted, I had every question I asked ( as you write an equation when you click on the equation box) in a word folder, but now they ar egone :/